# An Intention-driven Lane Change Framework Considering Heterogeneous Dynamic Cooperation in Mixed-traffic Environment

### v2

@@ Abstract
keywords:
mixed-traffic environment , lane change decision-making , human driving style , inverse reinforcement learning , intention-driven interaction , NGSIM
journal: Transportation Research Part C: Emerging Technologies
1 Introduction
The rapid advancement of autonomous driving technology has accelerated the transition toward mixed-traffic environments where autonomous vehicles (AVs) will inevitably share roadways with human-driven vehicles (HVs) for the foreseeable future (Yurtsever et al. (2020)). Within this operational paradigm, the ability of AVs to accurately interpret and adapt to human drivers’ heterogeneous behavioral patterns during lane change maneuvers emerges as a critical safety imperative (Xing et al. (2021); Chen et al. (2022)). This challenge is particularly acute given the fundamental uncertainty inherent in human decision-making processes and the dynamic complexity of traffic interactions (Schwarting et al. (2018)).
Modern lane change algorithms have evolved from classical models (Gipps (1986)) to contemporary data-driven approaches (Zheng (2014); Wang et al. (2019b)), and can be broadly categorized into three paradigms: rule-based, data-driven, and incentive-based systems. Recent advances in inverse reinforcement learning (Sun et al. (2022)) and probabilistic modeling (Sun et al. (2018)) have partially incorporated driver intentions into decision-making (Sheng et al. (2022)). Nevertheless, several critical limitations persist. Firstly, existing frameworks often rely on a behavioral homogeneity assumption, which oversimplifies surrounding human drivers by adopting uniform behavioral models and thereby overlooks inter-driver heterogeneity in decision-making. Secondly, although some methods incorporate personalized intentions, they typically fail to capture the temporal evolution of interactions among surrounding vehicles. Lastly, there remains an intention modeling gap: current prediction approaches lack mechanisms to effectively link observable driving styles with latent reward structures, resulting in a disconnect between behavior recognition and motion-planning.
These limitations are particularly problematic given that sudden lane changes account for about 17.0% of total severe crashes (Shawky (2020)), underscoring the urgent need for AV frameworks capable of simultaneously predicting trajectories and adapting to individual driving styles (Kuderer et al. (2015)).
To address the identified shortcomings in existing algorithms, we introduce an intention-driven lane change framework that synergistically integrates driving style recognition with adaptive decision-making as shown in Fig.1. In the context of this research, “driving style” is conceptualized as a habitual manner of vehicle operation, distinctive of an individual or subset of drivers, that serves as a proxy for underlying personalized driving intentions. Our proposed lane change framework incorporates driving style into two distinct yet interrelated components: intention-prediction and motion-planning. The main contributions of this study are summarized as follows:
  • 1.
@@ 1 Introduction
2 Related Work
2.1 Automated Lane Change Algorithms
Lane change algorithms have evolved from Gipps’ seminal car-following model (Gipps (1986)) into three methodological paradigms: rule-based, data-driven, and incentive-based approaches (Zheng (2014); Wang et al. (2019b)). These paradigms reflect the shift from handcrafted heuristics toward learning- and optimization-driven strategies, yet each exhibits inherent strengths and weaknesses.
Rule-based methods rely on predefined criteria to reproduce lane change behaviors. Representative studies include MOBIL (Kesting et al. (2007)), which minimizes overall braking induced by lane changes, and fuzzy-rule systems combined with LSTMs for maneuver forecasting (Wang et al. (2021)). Rule-based constraints have also been integrated into longitudinal control (Wang et al. (2019a)) and optimization formulations (He et al. (2021)). Although interpretable and computationally efficient, such approaches are rigid and sensitive to rule specification, often resulting in ambiguous decisions under complex traffic conditions.
Data-driven methods exploit large-scale trajectory datasets to learn nonlinear mappings between traffic states and lane change decisions. Classical examples include SVM-based classifiers (Mandalia and Salvucci (2005)) and ANN–SVM hybrids (Dou et al. (2016)), while deep learning has enabled richer feature representation through DNNs (Liu et al. (2019a)). Recent work further leverages Transformer architectures to capture temporal dependencies (Guo et al. (2025)). Despite improved accuracy, data-driven models often lack interpretability and generalizability, making their deployment in heterogeneous traffic environments non-trivial.
Incentive-based methods formulate lane change behavior as a sequential decision-making problem guided by rewards. Reinforcement learning has demonstrated promising results in adaptive driving policies (Hoel et al. (2019); Zhu et al. (2018b)), but suffers from reward design difficulties and sparse exploration. Inverse reinforcement learning (IRL) alleviates this by recovering latent reward structures from demonstrations (Arora and Doshi (2021); Sun et al. (2022); Huang et al. (2021)), enabling more human-like behavior synthesis. Nevertheless, IRL approaches remain computationally demanding and limited by state-space coverage.
@@ 2.1 Automated Lane Change Algorithms
2.2 Lane Change Frameworks Considering Intention Prediction
To enhance safety and efficiency, recent lane change frameworks increasingly incorporate the prediction of surrounding HVs intentions. A number of studies have integrated intention prediction into decision-making pipelines (Sun et al. (2022, 2018); Sheng et al. (2022)).
Sun et al. (2018) proposed a hierarchical decision-making model in which the first layer interacts with nearby vehicles to generate preliminary lane change decisions, while the second layer refines these into detailed trajectories. Guo et al. (2019) improved adaptive cruise control by embedding merging-behavior prediction into car-following dynamics, thereby enhancing comfort and safety. Leveraging deep sequence models, Gao et al. (2023) developed a dual-Transformer architecture that jointly predicts surrounding HVs’ intentions and trajectories, demonstrating high accuracy by explicitly modeling social interactions. Zhao et al. (2023) emphasized intention prediction within responsibility-sensitive safety frameworks, highlighting its role in proactive risk mitigation. More recently, Wang et al. (2025) introduced a socially aware DAG-based intention-prediction algorithm, which further improves reliability by capturing structural dependencies among surrounding vehicles.
@@ 2.2 Lane Change Frameworks Considering Intention Prediction
2.3 Human Personality Integrated Lane Change Frameworks
In mixed-traffic environments, frameworks that integrate human personality factors offer more adaptive and realistic decision-making for AVs. Human factors have therefore been incorporated into both decision-making and motion-planning modules. Sheng et al. (2022) designed a human-factor evaluation index emphasizing safety, efficiency, and comfort, which was subsequently embedded into a motion-planning algorithm that coordinated AV decisions with the interactive behaviors of surrounding vehicles. Liu et al. (2019b) incorporated benefit, safety, and tolerance factors into a support vector machine with Bayesian optimization to improve lane change rationality. Personalized driver traits have also been modeled explicitly: Butakov and Ioannou (2014), Liu et al. (2022), and Rosbach et al. (2019) developed strategies that adapt lane change policies to driver-specific characteristics. Zhu et al. (2018a) proposed a personalized assistance system that prioritizes safety by classifying drivers through a backpropagation neural network optimized with particle swarm methods. More recently, Liu and Peeta (2025) introduced a human-like strategy that considers AV–HV interactions, thereby improving performance in mixed-traffic settings.
@@ 3.1 Problem Formulation
To mimic human driving processes, we model the environment as a Markov decision process (MDP) with state $S$, action $A$, and reward $R$. The state space $S$ comprises both ego-vehicle states (velocity $v_{E}$, acceleration $a_{E}$) and surrounding-vehicle observations, including the velocities of the leading and target-rear vehicles $(v_{f},v_{tr})$ and their relative distances to the ego vehicle $(d_{Ef},d_{Etr})$. These features are grouped into inner features (ego-centric) and inter features (interaction-related). The action space $A=\{\textit{LC},\textit{LK}\}$ corresponds to either initiating a lane change or maintaining the current lane. The reward $R$ represents the cumulative return defined by efficiency, safety, and comfort, and will be formally specified in Section 4.
@@
3.2 Dataset
This study employs the Next Generation SIMulation (NGSIM) I-80 dataset (Halkia and Colyar (2006)) for both training and evaluation. The dataset was collected on a section of the Interstate 80 freeway in California using multiple synchronized video cameras installed on nearby high-rise buildings. The recorded videos were subsequently processed to extract vehicle trajectories at a sampling rate of 10 Hz. The dataset provides detailed microscopic information such as vehicle position, velocity, and lane assignment, offering a naturalistic and high-resolution representation of mixed-traffic conditions. Owing to its comprehensiveness and accuracy, the NGSIM I-80 dataset has become one of the most widely adopted benchmarks for studying lane change behaviors and human–vehicle interactions.
@@ 15 $t\leftarrow t+\text{dir}$
16
17 end while
18return 0 if $\text{dir}=-1$ else $\text{len}(v)-1$, False
@@
Figure 3: Performance of the lane change start and end point detection algorithm. The figure depicts the spatial distribution of lane change points along vehicle trajectories, with red dots representing the start points and green dots indicating the end points. The lane change duration, measured from the start to the end point, reflects the effectiveness and accuracy of the algorithm.
3.3 Lane Change Start-End Point Detection
@@
Upon identifying the start and end points, we compute the lane change duration. This methodology is applied to the lane change trajectories extracted from our dataset, enabling the determination of their respective lane change durations. As shown in Fig. 2, the log-normal fit distribution of lane change duration times, along with its residuals plot, aligns closely with findings from previous research (Li et al. (2023); Toledo and Zohar (2007)), substantiating the precision of our algorithms in accurately identifying lane change start and end points.
4 Methodology
@@
4.1 Driving Style Clustering and Recognition Module
@@
Table 1: Features selected for driving style clustering
@@
From each vehicle trajectory, six statistical features are computed, including mean and maximum velocity/acceleration and their standard deviations (see Table 1). To alleviate dimensionality issues, features are transformed via principal component analysis (PCA). K-means clustering is then applied to group drivers into three canonical styles (aggressive, normal, conservative), by minimizing the intra-cluster squared Euclidean distance. This unsupervised step provides discrete style labels without requiring manual annotation.
@@ 21     $\quad+\,\lambda_{2}\frac{1}{B}\sum_{i}\Big[(r^{(a_{i})}_{i})^{2}+(r^{(1-a_{i})}_{i})^{2}\Big]$
4.2 Intention-prediction Module
@@
4.3 Decision-making Module
@@
4.4 Motion-planning Module
@@
To predict the motion of surrounding vehicles, we adopt maximum-entropy inverse reinforcement learning (Max-Ent IRL) (Ziebart et al. (2008)). Unlike behavior cloning, which directly maps states to actions, IRL infers the latent reward function $R(s,a)$ that best explains expert demonstrations. Under the Max-Ent principle, the probability of a trajectory $\zeta$ is
@@
Following established formulations (Qiu et al. (2024)), the reward function is modeled as a linear combination of interpretable driving features:
@@
where $\Phi(s,a)$ encodes three aspects of human driving behavior:
  • 1.
    Efficiency:
@@
    reflecting the vehicle’s velocity magnitude.
  • 2.
    Safety:
@@
    where $x_{lv}$ and $x_{ego}$ denote the leading and ego vehicle positions, respectively, and $\epsilon$ avoids division by zero. This penalizes short headways.
  • 3.
    Comfort:
@@
    where $\dot{a}_{ego}(t)$ is the jerk (derivative of acceleration), discouraging abrupt maneuvers.
@@
With the lane change command of the proposed decision-making model, the AV is able to generate a lane change path based on the predicted trajectory of the target vehicle. Then MPC is used to control the AV drive from the current lane to the target lane considering the constraint of collision avoidance with the surrounding vehicles. Given the lane change command and the predicted trajectory of the following vehicle in the target lane, we generate a smooth lane change trajectory based on the Sigmoid function. To sufficiently utilize the predicted trajectory of the target vehicle, we generate the reference path of the AV using the Sigmoid function inspired by Ammour et al. (2020). With the input of the predicted trajectory of the target vehicle, and the current and target state of the ego vehicle, the expression of the sigmoid function-based trajectory is elaborated as follows. The sigmoid function-based trajectory incorporates the predicted trajectory of the target vehicle along with the current and target states of the ego vehicle. With the expected longitudinal distance $d_{\text{long}}$, lateral distance $d_{\text{lat}}$ during the lane change process, and the value of a curvature-related variable $\tau$, the reference path is generated by:
@@
The generated reference trajectory is then integrated into the MPC controller for trajectory tracking and safety guarantee. We use a bicycle model to simulate the dynamics of the ego vehicle. Afterward, MPC is used to track the reference trajectory for lane change maneuvers. Next, following the recipe proposed in Liu et al. (2025), each line of the road and the surrounding vehicles are modeled using potential functions integrated into the MPC-based trajectory.
5 Experiments and Results
5.1 Hyperparameter and Data Settings
@@
Figure 4: Analysis of driving style representation: (a) clustering results show distinct style separation; (b) feature correlation matrix highlights inter-feature dependencies.
5.2 Main Results
  • 1.
    IDM+MOBIL: A rule-based baseline that integrates the Intelligent Driver Model (IDM) for longitudinal control (Kesting et al. (2010)) with the MOBIL framework for lateral maneuvers (Kesting et al. (2007)). Lane change decisions are triggered by predefined safety constraints and acceleration incentives, aiming to balance efficiency and safety.
  • 2.
    H-LSTM: A learning-based approach employing a hierarchical long short-term memory (H-LSTM) network (Liao et al. (2022)) to capture temporal dependencies in multi-vehicle interactions and enable online lane change prediction.
  • 3.
    BC: A supervised imitation-learning baseline (Wang et al. (2023)), where the policy directly maps state features to expert actions via behavior cloning, without explicit reward modeling.
  • 4.
    VWC: A learning-based baseline employing a variable-weight cooperation (VWC) model (Wang et al. (2024)), which adaptively assigns dynamic weights to lane change conditions and driving demands across different phases of interaction.
Evaluation metrics for lane change intention prediction include precision, recall, and the F1 score, defined as follows (Frossard et al. (2019)):
precision $=\frac{TP}{TP+FP}$ (20)
recall $=\frac{TP}{TP+FN}$ (21)
Table 2 compares the performance of different baselines against our proposed method on LK and LC classification. The results clearly show that rule-based IDM+MOBIL performs the worst, with overall accuracy of only 0.6185, reflecting its limited adaptability to heterogeneous human driving patterns. Learning-based methods achieve significantly higher performance: H-LSTM reaches above 0.91 across all metrics by capturing temporal dependencies, while BC attains moderate performance (accuracy 0.8461) but is constrained by its reliance on direct imitation. The VWC model further improves classification accuracy to 0.9207 by adaptively weighting interactive features. In contrast, our proposed intention-driven BC-IRL framework achieves the best overall results, with accuracy 0.9418 and F1-score 0.9427. Notably, it improves lane change recognition (precision 0.8911, recall 0.9239, F1-score 0.9072) over all baselines, highlighting its ability to model inter-driver heterogeneity through cooperation scores and reward inference.
Table 2: Comparison of different methods on lane keeping (LK) and lane change (LC) classification.

### v3

@@ Abstract
I Introduction
The rapid advancement of autonomous driving technology has accelerated the transition toward mixed-traffic environments where autonomous vehicles (AVs) will inevitably share roadways with human-driven vehicles (HVs) for the foreseeable future [1]. Within this operational paradigm, the ability of AVs to accurately interpret and adapt to human drivers’ heterogeneous behavioral patterns during lane change maneuvers emerges as a critical safety imperative [2, 3]. This challenge is particularly acute given the fundamental uncertainty inherent in human decision-making processes and the dynamic complexity of traffic interactions [4].
Modern lane change algorithms have evolved from classical models [5] to contemporary data-driven approaches [6, 7], and can be broadly categorized into three paradigms: rule-based, data-driven, and incentive-based systems. Recent advances in inverse reinforcement learning [8] and probabilistic modeling [9] have partially incorporated driver intentions into decision-making [10]. Nevertheless, several critical limitations persist. Firstly, existing frameworks often rely on a behavioral homogeneity assumption, which oversimplifies surrounding human drivers by adopting uniform behavioral models and thereby overlooks inter-driver heterogeneity in decision-making. Secondly, although some methods incorporate personalized intentions, they typically fail to capture the temporal evolution of interactions among surrounding vehicles. Lastly, there remains an intention modeling gap: current prediction approaches lack mechanisms to effectively link observable driving styles with latent reward structures, resulting in a disconnect between behavior recognition and motion-planning.
These limitations are particularly problematic given that sudden lane changes account for about 17.0% of total severe crashes [11], underscoring the urgent need for AV frameworks capable of simultaneously predicting trajectories and adapting to individual driving styles [12].
To address the identified shortcomings in existing algorithms, we introduce an intention-driven lane change framework that synergistically integrates driving-style recognition with adaptive decision-making as shown in Fig.1. In the context of this research, “driving style” is conceptualized as a habitual manner of vehicle operation, distinctive of an individual or subset of drivers, that serves as a proxy for underlying personalized driving intentions. Our proposed lane change framework incorporates driving style into two distinct yet interrelated components: intention-prediction and motion-planning. The main contributions of this study are summarized as follows:
@@ I Introduction
II Related Work
II-A Automated Lane Change Algorithms
Lane change algorithms have evolved from Gipps’ seminal car-following model [5] into three methodological paradigms: rule-based, data-driven, and incentive-based approaches [6, 7]. These paradigms reflect the shift from handcrafted heuristics toward learning- and optimization-driven strategies, yet each exhibits inherent strengths and weaknesses.
Rule-based methods rely on predefined criteria to reproduce lane change behaviors. Representative studies include MOBIL [13], which minimizes overall braking induced by lane changes, and fuzzy-rule systems combined with LSTMs for maneuver forecasting [14]. Rule-based constraints have also been integrated into longitudinal control [15] and optimization formulations [16]. Although interpretable and computationally efficient, such approaches are rigid and sensitive to rule specification, often resulting in ambiguous decisions under complex traffic conditions.
Data-driven methods exploit large-scale trajectory datasets to learn nonlinear mappings between traffic states and lane change decisions. Classical examples include SVM-based classifiers [17] and ANN–SVM hybrids [18], while deep learning has enabled richer feature representation through DNNs [19]. Recent work further leverages Transformer architectures to capture temporal dependencies [20]. Despite improved accuracy, data-driven models often lack interpretability and generalizability, making their deployment in heterogeneous traffic environments non-trivial.
Incentive-based methods formulate lane change behavior as a sequential decision-making problem guided by rewards. Reinforcement learning has demonstrated promising results in adaptive driving policies [21, 22], but suffers from reward design difficulties and sparse exploration. Inverse reinforcement learning (IRL) alleviates this by recovering latent reward structures from demonstrations [23, 8, 24], enabling more human-like behavior synthesis. Nevertheless, IRL approaches remain computationally demanding and limited by state-space coverage.
@@ II-A Automated Lane Change Algorithms
II-B Lane Change Frameworks Considering Intention Prediction
To enhance safety and efficiency, recent lane change frameworks increasingly incorporate the prediction of surrounding HVs intentions. A number of studies have integrated intention prediction into decision-making pipelines [8, 9, 10].
Sun et al. [9] proposed a hierarchical decision-making model in which the first layer interacts with nearby vehicles to generate preliminary lane change decisions, while the second layer refines these into detailed trajectories. Guo et al. [25] improved adaptive cruise control by embedding merging-behavior prediction into car-following dynamics, thereby enhancing comfort and safety. Leveraging deep sequence models, Gao et al. [26] developed a dual-Transformer architecture that jointly predicts surrounding HVs’ intentions and trajectories, demonstrating high accuracy by explicitly modeling social interactions. Zhao et al. [27] emphasized intention prediction within responsibility-sensitive safety frameworks, highlighting its role in proactive risk mitigation. More recently, Wang et al. [28] introduced a socially aware DAG-based intention-prediction algorithm, which further improves reliability by capturing structural dependencies among surrounding vehicles.
@@ II-B Lane Change Frameworks Considering Intention Prediction
II-C Human Personality Integrated Lane Change Frameworks
In mixed-traffic environments, frameworks that integrate human personality factors offer more adaptive and realistic decision-making for AVs. Human factors have therefore been incorporated into both decision-making and motion-planning modules. Sheng et al. [10] designed a human-factor evaluation index emphasizing safety, efficiency, and comfort, which was subsequently embedded into a motion-planning algorithm that coordinated AV decisions with the interactive behaviors of surrounding vehicles. Liu et al. [29] incorporated benefit, safety, and tolerance factors into a support vector machine with Bayesian optimization to improve lane change rationality. Personalized driver traits have also been modeled explicitly: [30], [31], and [32] developed strategies that adapt lane change policies to driver-specific characteristics. Zhu et al. [33] proposed a personalized assistance system that prioritizes safety by classifying drivers through a backpropagation neural network optimized with particle swarm methods. More recently, Liu et al. [34] introduced a human-like strategy that considers AV–HV interactions, thereby improving performance in mixed-traffic settings.
@@ III-A Problem Formulation
To mimic human driving processes, we model the environment as a Markov decision process (MDP) with state $S$, action $A$, and reward $R$. The state space $S$ comprises both ego-vehicle states (velocity $v_{E}$, acceleration $a_{E}$) and surrounding-vehicle observations, including the velocities of the leading and target-rear vehicles $(v_{f},v_{tr})$ and their relative distances to the ego vehicle $(d_{Ef},d_{Etr})$. These features are grouped into inner features (ego-centric) and inter features (interaction-related). The action space $A=\{\textit{LC},\textit{LK}\}$ corresponds to either initiating a lane change or maintaining the current lane. The reward $R$ represents the cumulative return defined by efficiency, safety, and comfort, and will be formally specified in Section IV.
@@
III-B Dataset
This study employs the Next Generation SIMulation (NGSIM) I-80 dataset [35] for both training and evaluation. The dataset was collected on a section of the Interstate 80 freeway in California using multiple synchronized video cameras installed on nearby high-rise buildings. The recorded videos were subsequently processed to extract vehicle trajectories at a sampling rate of 10 Hz. The dataset provides detailed microscopic information such as vehicle position, velocity, and lane assignment, offering a naturalistic and high-resolution representation of mixed-traffic conditions. Owing to its comprehensiveness and accuracy, the NGSIM I-80 dataset has become one of the most widely adopted benchmarks for studying lane change behaviors and human–vehicle interactions.
@@ 14 $t\leftarrow t+\text{dir}$
15
16return 0 if $\text{dir}=-1$ else $\text{len}(v)-1$, False
@@
Figure 3: Performance of the lane change start and end point detection algorithm. Spatial distribution of lane change points along vehicle trajectories is illustrated. Red dots denote the initiation points and green dots the termination points. The lane change duration, measured from the start to the end point, reflects the effectiveness and accuracy of the algorithm.
III-C Lane Change Start-End Point Detection
@@
Upon identifying the start and end points, we compute the lane change duration. This methodology is applied to the lane change trajectories extracted from our dataset, enabling the determination of their respective lane change durations. As shown in Fig. 2, the log-normal fit distribution of lane change duration times, along with its residuals plot, aligns closely with findings from previous research [36, 37], substantiating the precision of our algorithms in accurately identifying lane change start and end points.
IV Methodology
@@
IV-A Driving Style Clustering and Recognition Module
@@
TABLE I: Features selected for driving style clustering
@@
From each vehicle trajectory, six statistical features are computed, including mean and maximum velocity/acceleration and their standard deviations (see Table I). To alleviate dimensionality issues, features are transformed via principal component analysis (PCA). K-means clustering is then applied to group drivers into three canonical styles (aggressive, normal, conservative), by minimizing the intra-cluster squared Euclidean distance. This unsupervised step provides discrete style labels without requiring manual annotation.
@@ 21     $\quad+\,\lambda_{2}\frac{1}{B}\sum_{i}\Big[(r^{(a_{i})}_{i})^{2}+(r^{(1-a_{i})}_{i})^{2}\Big]$
IV-B Intention-prediction Module
@@
IV-C Decision-making Module
@@
IV-D Motion-planning Module
@@
To predict the motion of surrounding vehicles, we adopt maximum-entropy inverse reinforcement learning (Max-Ent IRL) [38]. Unlike behavior cloning, which directly maps states to actions, IRL infers the latent reward function $R(s,a)$ that best explains expert demonstrations. Under the Max-Ent principle, the probability of a trajectory $\zeta$ is
@@
Following established formulations [39], the reward function is modeled as a linear combination of interpretable driving features:
@@
where the feature vector $\Phi(s,a)$ encodes three fundamental aspects of human driving behavior: Efficiency, Safety, and Comfort.
The Efficiency feature, $f_{\text{eff}}(t)$, reflects the vehicle’s velocity magnitude, defined as:
@@
The Safety feature, $f_{\text{safety}}(t)$, is designed to penalize short headways, ensuring a safe following distance, formulated as:
@@
where $x_{lv}$ and $x_{ego}$ denote the positions of the leading and ego vehicles, respectively, and $\epsilon$ is a small constant introduced to prevent division by zero.
The Comfort feature, $f_{\text{comfort}}(t)$, discourages abrupt maneuvers by incorporating the jerk (the derivative of acceleration, $\dot{a}_{ego}(t)$) into its definition:
@@
where $\dot{a}_{ego}(t)$ is the jerk of the ego vehicle at time $t$.
@@
With the lane change command of the proposed decision-making model, the AV is able to generate a lane change path based on the predicted trajectory of the target vehicle. Then MPC is used to control the AV drive from the current lane to the target lane considering the constraint of collision avoidance with the surrounding vehicles. Given the lane change command and the predicted trajectory of the following vehicle in the target lane, we generate a smooth lane change trajectory based on the Sigmoid function. To sufficiently utilize the predicted trajectory of the target vehicle, we generate the reference path of the AV using the Sigmoid function inspired by [40]. With the input of the predicted trajectory of the target vehicle, and the current and target state of the ego vehicle, the expression of the sigmoid function-based trajectory is elaborated as follows. The sigmoid function-based trajectory incorporates the predicted trajectory of the target vehicle along with the current and target states of the ego vehicle. With the expected longitudinal distance $d_{\text{long}}$, lateral distance $d_{\text{lat}}$ during the lane change process, and the value of a curvature-related variable $\tau$, the reference path is generated by:
@@
The generated reference trajectory is then integrated into the MPC controller for trajectory tracking and safety guarantee. We use a bicycle model to simulate the dynamics of the ego vehicle. Afterward, MPC is used to track the reference trajectory for lane change maneuvers. Next, following the recipe proposed in [41], each line of the road and the surrounding vehicles are modeled using potential functions integrated into the MPC-based trajectory.
@@
Figure 4: Analysis of driving-style representation: (a) clustering results show distinct style separation; (b) feature correlation matrix highlights inter-feature dependencies.
V Experiments and Results
V-A Hyperparameter and Data Settings
TABLE II: Comparison of different methods on lane keeping (LK) and lane change (LC) classification.